# Mitigating Propensity Bias of Large Language Models for Recommender Systems

### v1

@@ 4.3. Counterfactual Debiased Recommendation
  • $T\rightarrow P$ and $S\rightarrow O$ indicate that the side information generated by the LLM is embedded with the LLM’s own propensities.
  • $T\rightarrow I$ and $S\rightarrow U$ indicate that the usage of side information to improve recommendation performance.
  • $P\rightarrow I$ and $S\rightarrow U$ denote that the LLM’s propensity bias affects recommendation results.
Under the designed causal DAG, the problem of addressing the propensity bias of LLM becomes how to estimate the indirect effect of $O$ and $P$ on $U$ and $I$. Following the causal effect introduced in Section 3.3, we first construct the counterfactual-world causal graph shown in Figure 5 (b). Gray nodes represent reference states, and half-shaded nodes represent that they are influenced by the reference state. The total effect (TE) of individual $j$ ($j$ may be a user or a item) can be calculated as:
@@
The indirect paths (i.e., $S\rightarrow Y\rightarrow U$ and $T\rightarrow P\rightarrow I$ ) are blocked by setting $O$ and $P$ to reference states, i.e., $Y=Y(S=\left.s^{*}\right)$ and $P=P(T=\left.t^{*}\right)$. After that, we can calculate the PNDE of individual $j$ as:
@@
According to Eq. (3), we can eliminate the effect of propensity bias on recommendation results by reducing PNDE in TE:
@@ 4.4. Training and Inference
(17) $\mathbf{e}_{u}=\sigma\left(f_{P}\left(\mathbf{s}_{u}\right)\right)*\mathbf{\hat{e}}_{u},$
@@
(19) $\mathcal{L}_{\text{a}}=-\mathbb{E}\log\left[\frac{f\left(\mathbf{e}_{u},\mathbf{s}_{u}\right)}{\sum_{u\in m}f\left(\mathbf{e}_{u},\mathbf{s}_{u}\right)}\right]-\mathbb{E}\log\left[\frac{f\left(\mathbf{e}_{i},\mathbf{s}_{i}\right)}{\sum_{i\in\mathbf{I}}f\left(\mathbf{e}_{i},\mathbf{s}_{i}\right)}\right],$
@@
where $O=\left\{\left(u,i^{+},i^{-}\right)\mid\left(u,i^{+}\right)\in\mathcal{R},\left(u,i^{-}\right)\in\mathcal{R}^{-}\right\}$ denotes the training set, and $\left(u,i^{+},i^{-}\right)$ is a training sample. Here, $\mathcal{R}$ is the set of observed user-item interactions, and $i^{+}$ is a positive sample that user $u$ has interacted with.
@@
(21) $\mathbf{\tilde{e}}_{u}=\mathbf{e}_{u}-\alpha*\sigma\left(f_{P}\left(\mathbf{s}_{u}\right)\right),$
@@ 5. Experimental evaluation
  • RQ4: How does each module in SSE function?
@@ 5.1.1. Datasets
Table 3. Instructions for obtaining side information using LLM on the three datasets.
@@
Amazon Item You will serve as an assistant to help me summarise which types of users would enjoy a specific book. I will provide you with the title and a description of the book. Here are the instructions: 1. I will provide you with information in the form of a JSON string that describes the book: “title”: “the title of the book”, “description”: “a description of the book”,
@@
User You will serve as an assistant to help me determine which types of books a specific user is likely to enjoy. I will provide you with information about books that the user has purchased, as well as his or her reviews of those books. Here are the instructions: 1. Each purchased book will be described in JSON format, with the following attributes: “title”: “the title of the book”, “description”: “a description of what types of users will like this book”, ”review”: “the user’s review on the book” 2. The information I will give you: PURCHASED ITEMS: a list of JSON strings describing the items that the user has purchased.
@@
Yelp Item You will serve as an assistant to help me summarise which types of users would enjoy a specific video game. I will provide you with the basic information (name, publisher, genres and tags) of that game and also some feedback of users for it. Here are the instructions: 1. The basic information will be described in JSON format, with the following attributes: “name”: “the name of the video game”, “publisher”: “the publisher of the game”, “genres”: “the genres of the game”, “tags”: “several tags describing the game” 2. Feedback from users will be managed in the following List format: [“the first feedback”, “the second feedback”, ….] 3. The information I will give you: BASIC INFORMATION: a JSON string describing the basic information about the game. USER FEEDBACK: a List object containing some feedback from users about the game.
@@
User You will serve as an assistant to help me determine which types of game a specific user is likely to enjoy. I will provide you with information about games that the user has interacted, as well as his or her reviews of those games. Here are the instructions: 1. Each interacted game will be described in JSON format, with the following attributes: “title”: “the name/title of the game”, “description”: “a description of what types of users will like this game”, “review”: “the user’s review on the game” 2. The information I will give you: PLAYED GAMES: a list of JSON strings describing the games that the user has played.
@@
Steam Item You will serve as an assistant to help me summarise which types of users would enjoy a specific business. I will provide you with the basic information (name, city and category) of that business and also some feedback of users for it. Here are the instructions: 1. The basic information will be described in JSON format, with the following attributes: “name”: “the name of the business”, “cit”: “city where the company is located”, (if there is no city, I will set this value to “None”) “categories”: “several tags describing the business” (if there is no categories, I will set this value to “None”) 2. Feedback from users will be managed in the following List format: [ “the first feedback”, “the second feedback”, …. ] 2. The information I will give you: BASIC INFORMATION: a JSON string describing the basic information about the business. USER FEEDBACK: a List object containing some feedback from users about the business.
@@
User You will serve as an assistant to help me determine which types of business a specific user is likely to enjoy. I will provide you with information about businesses that the user has interacted, as well as his or her reviews of those businesses. Here are the instructions: 1. Each interacted business will be described in JSON format, with the following attributes: “title”: “the name of the business”, “description”: “a description of what types of users will like this business”, “review”: “the user’s review on the business” 2. The information I will give you: PURCHASED BUSINESSES: a list of JSON strings describing the businesses that the user has interacted.
@@
Table 4. Comparison of recommendation performance of all methods on different datasets, where improvement is abbreviated as “Imprv.”.
@@ 5.1.2. Evaluation Metrics
To ensure a thorough and comprehensive evaluation of our proposed CLLMR method, we employed a ranking scheme across all experiments, consistent with methodologies from previous studies (He et al., 2020; Ren et al., 2024). We selected two well-established ranking-based metrics to quantify the effectiveness of our approach: Recall@N and NDCG@N. By using these metrics, we aim to demonstrate the ability of our method to deliver high-quality results that align with user expectations and preferences.
@@ 5.1.4. Baselines
5.2. Performance Comparison
5.3. Ablation Study
5.3.1. Necessity of each module of CLLMR
@@
Table 5. The experimental results of ablation studies on three datasets. Removing the component of counterfactual inference is abbreviated as “w/o Conf.”, and removing the component of the spectrum-based side information encoder is abbreviated as “w/o SSE”.
@@ 5.4. Hyperparameter Analysis
In this section, we conduct a hyperparameter analysis for the rank of the spectrum in SSE to answer RQ5, with the experimental results shown in Figure 10. We set the rank to 16, 32, and 64, respectively, and it is important to note that the hidden layer dimension is set to 32 for all the experiments in this paper. The results show that our proposed method achieves the best performance when the rank is set to 32.

### v2

@@ 4.3. Counterfactual Debiased Recommendation
  • $T\rightarrow P$ and $S\rightarrow O$ denote that part of the side representation is affected by the user/item propensity bias.
  • $T\rightarrow I$ and $S\rightarrow U$ denote that the side representation is affected by both the user/item unrelated and user/item propensity biases, which directly affects the user/item collaborative representation.
  • $P\rightarrow I$ and $O\rightarrow U$ denote that user/item side representations affected by the user/item propensity bias are aligned with user/item collaborative representations to improve recommendation performance.
The aim of this paper is to mitigate the effect of user/item unrelated propensity bias on recommendations. Under the designed causal DAG, the problem of mitigating the unrelated propensity bias of the LLM becomes how to estimate the indirect effects of $O$ and $P$ on $U$ and $I$. Following the causal effect introduced in Section 3.3, we first construct the counterfactual-world causal graph shown in Figure 5 (b). Gray nodes represent reference states, and half-shaded nodes represent that they are influenced by the reference state. As we discussed in Section 3.3, to ensure a valid causal interpretation of the PNDE and TNIE in the causal graph in Figure 5, the following assumptions must hold:
  • No unmeasured confounding of the exposure-outcome effect: All confounders that affect both the exposure $S$ or $T$ and the outcome $U$ or $I$ must be observed and appropriately adjusted for. This ensures that the direct effects are not confounded by unmeasured variables.
  • No unmeasured confounding of the mediator-outcome effect: All variables influencing the mediator $O$ or $P$ and the outcome $U$ or $I$ must be measured. This ensures that the indirect effects are not biased by hidden confounders.
  • No unmeasured confounding of the exposure-mediator effect: Every variable that confounds the relationship between the exposure $S$ or $T$ and the mediator $O$ or $P$ must be observed and accounted for. This ensures that the causal effects are not confounded by unmeasured variables.
  • No confounders of the mediator-outcome effect that are affected by the exposure: This assumption requires that no variables confounding the mediator-outcome relationship $O\rightarrow U$ or $P\rightarrow I$ are themselves influenced by the exposure $S$ or $T$.
These assumptions ensure that the causal effects are identified without bias from unmeasured variables or violations of the causal structure implied by the graph. The total effect (TE) of individual $j$ ($j$ may be a user or a item) can be calculated as:
@@
The indirect paths (i.e., $S\rightarrow O\rightarrow U$ and $T\rightarrow P\rightarrow I$ ) are blocked by setting $O$ and $P$ to reference states, i.e., $Y=Y(S=\left.s^{*}\right)$ and $P=P(T=\left.t^{*}\right)$. After that, we can calculate the PNDE of individual $j$ as:
@@
According to Eq. (3), we can eliminate the effect of unrelated propensity bias on recommendation results by reducing PNDE in TE:
@@ 4.4. Training and Inference
(17) $\mathbf{e}_{u}=\sigma\left(f_{O}\left(\mathbf{s}_{u}\right)\right)*\mathbf{\hat{e}}_{u},$
@@
(19) $\mathcal{L}_{\text{a}}=-\mathbb{E}\log\left[\frac{f\left(\mathbf{e}_{u},\mathbf{s}_{u}\right)}{\sum_{u\in\mathbf{U}}f\left(\mathbf{e}_{u},\mathbf{s}_{u}\right)}\right]-\mathbb{E}\log\left[\frac{f\left(\mathbf{e}_{i},\mathbf{s}_{i}\right)}{\sum_{i\in\mathbf{I}}f\left(\mathbf{e}_{i},\mathbf{s}_{i}\right)}\right],$
@@
where $O=\left\{\left(u,i^{+},i^{-}\right)\mid\left(u,i^{+}\right)\in\mathcal{R},\left(u,i^{-}\right)\in\mathcal{R}^{-}\right\}$ denotes the training set, and $\left(u,i^{+},i^{-}\right)$ is a training sample. Specifically, $i^{+}$ denotes an item that has interacted with $u$ and $i^{-}$ denotes an item that has not interacted with $u$. Here, $\mathcal{R}$ is the set of observed user-item interactions, and $i^{+}$ is a positive sample that user $u$ has interacted with.
@@
(21) $\mathbf{\tilde{e}}_{u}=\mathbf{e}_{u}-\alpha*\sigma\left(f_{O}\left(\mathbf{s}_{u}\right)\right),$
@@ 5. Experimental evaluation
  • RQ4: Is each module in SSE function necessary?
@@ 5.1.1. Datasets
Table 3. Instructions for obtaining side information using LLM on the three datasets following (Ren et al., 2024). “You … summarise” is short for “You will serve as an assistant to help me summarise”. “You … determine” is short for “You will serve as an assistant to help me determine”.
@@
Amazon Item You … summarise which types of users would enjoy a specific book. I will provide you with the title and a description of the book. Here are the instructions: 1. I will provide you with information in the form of a JSON string that describes the book: “title”: “the title of the book”, “description”: “a description of the book”,
@@
User You … determine which types of books a specific user is likely to enjoy. I will provide you with information about books that the user has purchased, as well as his or her reviews of those books. Here are the instructions: 1. Each purchased book will be described in JSON format, with the following attributes: “title”: “the title of the book”, “description”: “a description of what types of users will like this book”, ”review”: “the user’s review on the book” 2. The information I will give you: a list of JSON strings describing the items that the user has purchased.
@@
Yelp Item You … summarise which types of users would enjoy a specific video game. I will provide you with the basic information (name, publisher, genres and tags) of that game and also some feedback of users for it. Here are the instructions: 1. The basic information will be described in JSON format, with the following attributes: “name”: “the name of the video game”, “publisher”: “the publisher of the game”, “genres”: “the genres of the game”, “tags”: “several tags describing the game” 2. Feedback from users will be managed in the following List format: [“the first feedback”, “the second feedback”, ….] 3. The information I will give you: a JSON string describing the basic information about the game; a List object containing some feedback from users about the game.
@@
User You … determine which types of game a specific user is likely to enjoy. I will provide you with information about games that the user has interacted, as well as his or her reviews of those games. Here are the instructions: 1. Each interacted game will be described in JSON format, with the following attributes: “title”: “the name/title of the game”, “description”: “a description of what types of users will like this game”, “review”: “the user’s review on the game” 2. The information I will give you: a list of JSON strings describing the games that the user has played.
@@
Steam Item You … summarise which types of users would enjoy a specific business. I will provide you with the basic information (name, city and category) of that business and also some feedback of users for it. Here are the instructions: 1. The basic information will be described in JSON format, with the following attributes: “name”: “the name of the business”, “city”: “city where the company is located”, (if there is no city, I will set this value to “None”) “categories”: “several tags describing the business” (if there is no categories, I will set this value to “None”) 2. Feedback from users will be managed in the following List format: [ “the first feedback”, “the second feedback”, …. ] 3. The information I will give you: a JSON string describing the basic information about the business. a list object containing some feedback from users about the business.
@@
User You … determine which types of business a specific user is likely to enjoy. I will provide you with information about businesses that the user has interacted, as well as his or her reviews of those businesses. Here are the instructions: 1. Each interacted business will be described in JSON format, with the following attributes: “title”: “the name of the business”, “description”: “a description of what types of users will like this business”, “review”: “the user’s review on the business” 2. The information I will give you: a list of JSON strings describing the businesses that the user has interacted.
@@
Table 4. Comparison of recommendation performance of all methods on different datasets. The best value is bolded, the runner-up value is underlined, and the improvement over the runner-up by the best performer is abbreviated as “Imprv.”.
@@ 5.1.2. Evaluation Metrics
To ensure a thorough and comprehensive evaluation of our proposed CLLMR method, we employed a ranking scheme across all experiments, consistent with methodologies from previous studies (He et al., 2020; Ren et al., 2024). We selected two well-established ranking-based metrics to quantify the effectiveness of our approach: Recall@N and NDCG@N. By using these metrics, we aim to demonstrate the ability of our method to deliver high-quality results that align with user expectations and preferences. For clarity on the effectiveness of the improvement, we will calculate the magnitude of the improvement using the formula $\frac{b-a}{a}\times 100\%$, where $a$ is the best value and $b$ is the runner-up value.
@@ 5.2. Performance Comparison
5.3. Ablation Study
5.3.1. Necessity of each module of CLLMR
Table 5. The experimental results of ablation studies on three datasets. In the “Variants” column, “w/o Conf.” indicates removing the component of counterfactual inference, “w/o SSE” indicates removing the component of the spectrum-based side information encoder, and “De.” indicates the magnitude due to the removal of a component.
@@ 5.4. Hyperparameter Analysis
In this section, we do the hyperparameter analyses of the rank of the spectrum and the noise size in SSE, respectively.
The rank of the spectrum in SSE
In this subsection, we conduct a hyperparameter analysis for the rank of the spectrum in SSE to answer RQ5, with the experimental results shown in Figure 10. We set the rank to 16, 32, and 64, respectively, and it is important to note that the hidden layer dimension is set to 32 for all the experiments in other sections. The results show that our proposed method achieves the best performance when the rank is set to 32.
@@
(a) LRGCCF
(b) LightGCN
(c) SGL
(d) DirectAU
(e) SimGCL
(f) LightGCL
Figure 11. Hyperparameter analysis for the noise size in SSE.
The noise size in SSE
To better answer RQ4, we conduct a hyperparameter analysis of the noise size in SSE, with the experimental results shown in Figure 11. We evaluate the rank settings to 0.001, 0.005, 0.01, 0.05 and 0.1, respectively. It is important to note that the hidden layer dimension is fixed at 0.01 for all the experiments in other sections. The results show that our proposed method achieves the best performance when the rank is set to 0.01.
The experimental results reveal that the performance of the model gradually improves when the noise size is increased from 0.001 to 0.01. This proves that adding noise to the spectral information enhances the generalisation ability of the model by enabling it to learn more diverse data during training. However, when the noise increases beyond 0.1, the model’s effectiveness appears to decrease, or in some scenarios, remains unchanged. This is because as the noise increases, the information in the spectrum itself is affected or even overwritten, which makes the model performance converge. At the same time, we can observe that when the noise obtains a smaller value, the model performs better than when the noise is larger, which once again verifies that a small amount of noise helps to improve the generalisation ability of the model.
@@ 6. Conclusion
ACKNOWLEDGMENTS
This work has been supported in part by the National Natural Science Foundation of China under Grant 71774159, the China Postdoctoral Science Foundation under Grant 2021T140707, the Jiangsu Postdoctoral Science Foundation under grant number 2021K565C, the Science and Technology Foundation of Xuzhou under Grant KC22047, Australian Research Council under Grant DP230101122, the Graduate Innovation Program of China University of Mining and Technology 2024WLKXJ183, the Fundamental Research Funds for the Central Universities 2024-10949, the Postgraduate Research & Practice Innovation Program of Jiangsu Province KYCX24_2781, Research Fund of Guangxi Key Lab of Multi-source Information Mining & Security (MIMS24-13), and Guangxi Key Laboratory of Big Data in Finance and Economics (Grant No. FEDOP2022A03).